# GDEPO: Group Dual-dynamic and Equal-right-advantage Policy Optimization with Enhanced Training Data Utilization for Sample-Constrained Reinforcement Learning

### v1

@@ Abstract
Automated Theorem Proving (ATP) represents a fundamental challenge in Artificial Intelligence (AI), requiring the construction of machine-verifiable proofs in formal languages such as Lean to evaluate AI reasoning capabilities. Reinforcement learning (RL), particularly the high-performance Group Relative Policy Optimization (GRPO) algorithm, has emerged as a mainstream approach for this task. However, in ATP scenarios, GRPO faces two critical issues: when composite rewards are used, its relative advantage estimation may conflict with the binary feedback from the formal verifier; meanwhile, its static sampling strategy may discard entire batches of data if no valid proof is found, resulting in zero contribution to model updates and significant data waste. To address these limitations, we propose Group Dual-dynamic and Equal-right-advantage Policy Optimization (GDEPO), a method incorporating three core mechanisms: 1) dynamic additional sampling, which resamples invalid batches until a valid proof is discovered; 2) equal-right advantage, decoupling the sign of the advantage function (based on correctness) from its magnitude (modulated by auxiliary rewards) to ensure stable and correct policy updates; and 3) dynamic additional iterations, applying extra gradient steps to initially failed but eventually successful samples to accelerate learning on challenging cases. Experiments conducted on three datasets of varying difficulty (MinF2F-test, MathOlympiadBench, PutnamBench) confirm the effectiveness of GDEPO, while ablation studies validate the necessity of its synergistic components. The proposed method enhances data utilization and optimization efficiency, offering a novel training paradigm for ATP.
@@ I Introduction
The rapid advancement of Large Language Models (LLMs) has opened new avenues for applying Artificial Intelligence (AI) to formal reasoning. Automated Theorem Proving (ATP), a core challenge in AI, requires systems to construct machine-verifiable mathematical proofs in formal languages such as Lean [3, 4, 31]. Recent progress in ATP on complex theorem-proving tasks stems from the strong prior knowledge and generative capabilities of LLMs, coupled with the availability of extensive training and validation datasets [34, 14, 24, 5, 9, 26]. Notably, AlphaGeometry [23] have demonstrated problem-solving performance approaching International Mathematical Olympiad (IMO) medalist levels, signaling a transition of LLM-driven ATP from theoretical exploration toward practical deployment.
@@
To address the aforementioned issues, we propose the Group Dual-dynamic and Equal-right-advantage Policy Optimization (GDEPO) method, whose workflow is illustrated in Fig. 1. This approach identifies and discards ineffective training data, then re-explores valid solutions. It ensures correct optimization direction under arbitrary reward weight combinations through the equal-right advantage mechanism. Furthermore, for initially invalid training instances that yield correct proofs after re-exploration, the method applies a dynamic number of additional backward passes to rapidly assimilate knowledge that is challenging yet solvable for the model.
@@ II-B Inadequate Adaptability of Policy Optimization Algorithms in Formal Theorem Proving
Inspired by informal LLMs in natural language processing, recent foundational work in formal theorem proving has adopted similar reinforcement learning-based policy optimization strategies. Current formal theorem provers primarily rely on policy optimization algorithms originally developed for informal LLM training, such as Proximal Policy Optimization (PPO) [21], Direct Preference Optimization (DPO) [17], and Group Relative Policy Optimization (GRPO) [22], which is widely employed by state-of-the-art (SOTA) systems [19, 25, 14]. While these algorithms have demonstrated effectiveness in their native domains—such as language alignment and complex reasoning—they exhibit significant adaptation limitations when transferred to the formal theorem proving setting, where feedback is precise and binary.
@@ II-B3 Limitations of the Group Relative Policy Optimization Paradigm
GRPO has become the core optimization algorithm in current SOTA provers—such as DeepSeek-Prover-V2 [19], Kimina-Prover [25], and Goedel-Prover-V2 [14]—due to its elimination of a separate reward model and computational efficiency. However, its fundamental mechanism, which relies on relative comparisons among samples within a group, inherently conflicts with the absolute correctness criterion demanded by formal verification. This tension manifests in two critical issues. First, when composite rewards are employed, GRPO’s advantage computation may assign negative advantages to otherwise correct proof trajectories due to suboptimal scores in other reward components—a mathematical possibility rigorously established in Section VI. To avoid this, practitioners often resort to binary rewards alone, thereby forfeiting the core capability of reinforcement learning to align large language model outputs with nuanced human preferences, such as proof conciseness or elegance. Second, GRPO exhibits low data utilization efficiency: in batches containing a high proportion of challenging samples, using a single binary reward frequently results in zero advantages for all trajectories within a group. Consequently, all sampled attempts that fail to produce a valid proof are collectively discarded, wasting valuable training data. To mitigate these risks, existing SOTA systems [19, 25, 14] adopt a conservative strategy—using only binary rewards and deliberately forgoing the potential benefits of composite rewards. Moreover, data batches comprising exclusively correct or incorrect samples are often discarded outright [32]. These compromises yield dual adverse consequences: (i) the precise and rich feedback from formal verifiers remains underutilized, and (ii) the already scarce high-quality training data suffers from further reduced effective usage, impeding the model’s ability to learn from challenging examples. Additionally, the use of a fixed number of optimization iterations treats all samples uniformly, ignoring inherent differences in their learning value. As a result, the current training framework leaves considerable room for more efficient exploitation of formal proof data.
@@
The overall architecture of the GDEPO framework and its comparison with conventional paradigms are illustrated in Fig. 2. The informal large language model reinforcement learning paradigm, depicted in Fig. 2a, relies on a trained reward model to generate scalar rewards. However, such rewards lack the precise binary signal required to guarantee formal correctness. In contrast, the standard formal automated theorem proving paradigm, shown in Fig. 2b, employs a Lean verifier as the ground-truth judge but inherits two rigid design choices from generic reinforcement learning: (i) a single-round sampling mechanism that struggles to explore low-probability correct solutions for complex problems, and (ii) a fixed number of iterations applied uniformly across all training samples, disregarding their varying learning utility. Moreover, the advantage estimation in this paradigm faces a dilemma: using a composite reward function often leads to misaligned gradient directions, whereas relying solely on the binary correctness reward results in poor data efficiency and weakens the ability of RL optimization to reflect desired output preferences.
@@ III Group Dual-dynamic and Equal-Right Advantage Policy Optimization
GDEPO was proposed to address the aforementioned limitations of standard reinforcement learning in formal theorem proving. The framework simultaneously improves data efficiency and optimization rationality.
@@ III-B Equal-Right Advantage
Equal-right advantage decouples the supervision signal: the sign of the advantage—i.e., the direction of policy update—is determined solely by the binary label from the verifier, while auxiliary rewards modulate only the magnitude of the advantage—i.e., the update strength. This mechanism guarantees that all correct proofs receive positive signals and all incorrect ones receive negative signals, preserving the authority of verification outcomes while enabling multi-objective reward shaping. Importantly, it yields correct policy update directions even when all $G$ trajectories for a query are either entirely correct or entirely incorrect.
@@ III-C Dynamic Additional Iterations
Although the dynamic additional sampling module enhances the exploration of low-probability correct solutions, it introduces a new issue of sample heterogeneity: proofs obtained directly in the initial sampling and those discovered only after repeated exploration possess markedly different instructional values for model optimization. The latter—requiring sustained search to succeed—deserve deeper reinforcement learning attention. Applying identical single-step gradient updates to both cases fails to exploit the richer learning potential of the latter.
To address this, dynamic additional iterations are proposed to allocate computational resources proportionally to each sample’s learning value. The principle is straightforward: easy queries (correct on the first attempt) and persistently unsolvable hard queries undergo only one update iteration, whereas medium-difficulty queries—those that yield a correct proof only after multiple sampling rounds—are subjected to $m$ iterations to accelerate optimization. Formally, for a query $q$, if its valid output set $\mathcal{G}_{q}^{*}$ contains at least one correct trajectory and was not produced in the first sampling round, then all $G$ trajectories for $q$ are used in $m$ consecutive policy updates. The overall objective function is expressed as:
@@ IV-A Dataset for Training
The experiments are conducted on the publicly available Lean 4 [4] formalized mathematical theorem dataset. To rigorously evaluate GDEPO’s core capability—learning from challenging samples—we construct a targeted high-difficulty training set. The data is sourced from FineLeanCorpus [16], a high-quality dataset comprising formally verified propositions with broad domain coverage and inherently high difficulty. To further isolate highly challenging instances, we use the pass@32 performance of the DeepSeek-Prover-V2-671B model as a filtering criterion, retaining only those samples for which no correct proof is found across 32 independent sampling attempts.
@@ IV-C Baselines
We compare GDEPO-Prover with three SOTA models representing leading paradigms in the field of automated theorem proving, ensuring a diverse evaluation across model architectures and optimization strategies.
@@ IV-C1 DeepSeek-Prover-V2
DeepSeek-Prover-V2 [19] adopts a recursive subgoal decomposition paradigm. Its training process integrates multiple stages: problem decomposition and formalization using DeepSeek-V3; recursive subgoal solving with a 7 billion parameter model; curriculum learning for task generation; reinforcement learning via GRPO algorithm, complemented by supervised fine-tuning and knowledge distillation.
@@ IV-C2 Kimina-Prover
Kimina-Prover [25] emphasizes training stability through KL-constrained reinforcement learning. It initializes via few-shot fine-tuning, featuring innovations such as automated formalization methods for dataset construction, specialized tokenized reasoning formats for aligning informal and formal reasoning, and techniques like format filtering and negative gradient dropping to enhance stability.
@@ IV-C3 Goedel-Prover-V2
Goedel-Prover-V2 [14] is centered around scaffolded data synthesis and iterative self-correction. Key components include scaffolded data synthesis generating progressively challenging tasks, verifier-guided self-correction leveraging Lean compiler feedback for iterative proof correction, model averaging to mitigate output diversity decline in later training stages by blending model checkpoints, and hybrid GRPO reinforcement learning incorporating dynamic sampling [32], combined with supervised fine-tuning for efficient theorem proving with smaller parameter sizes.
These benchmark models cover three core technical directions in automated theorem proving—modular decomposition, constraint optimization, and feedback-loop enabled synthetic data. By comparing GDEPO, which focuses on innovative training algorithms, against these models, we can precisely isolate the performance gains attributed to our proposed dynamic sampling and advantage estimation mechanisms.
@@ IV-D Training Details
To ensure a fair comparison, all models share standardized common hyperparameters, while their model-specific parameters are individually tuned to best match the performance reported by the original authors. GDEPO is built upon the Godel-Prover-V2-8B base model [14] and extends the open-r1 algorithmic framework [10].
@@ V-A Overall Performance Evaluation
To evaluate the overall performance of GDEPO, we trained GOEDEL-Prover-V2 on a dataset of 30,000 samples containing 30% challenging instances and selected the best-performing model as the representative of our method. Table I reports the performance of this model alongside state-of-the-art baselines, including DeepSeek-Prover-V2 [19], Kimina-Prover [25], and Goedel-Prover-V2 [14], along with their variants of different parameter scales. The results show that GDEPO consistently improves performance across all three benchmarks, with a notable 84% gain on the most challenging PutnamBench. This demonstrates that GDEPO can effectively unlock additional performance potential even in models that have approached their performance ceilings, highlighting its strong capability in learning from challenging samples.
TABLE I: Performance Comparison of GDEPO-Prover with Existing State-of-the-Art Models and Their Variants Across Model Sizes.
@@
V-B Utilization Effect of Mislabeled Training Dataset Evaluation
@@ V-D Ablation Evaluation
TABLE III: Ablation study.
@@ VI-A Advantage Sign as Determinant of Trajectory Probability Direction
Here, $\pi_{\theta_{\text{old}}^{(s)}}(o_{i,t}^{*}\mid q,o_{i,<t}^{*})$ denotes the probability under the old policy and remains constant with respect to $\theta$. The expectation operator $\mathbb{E}$ is representing expectation over the question distribution and the effective response groups sampled from the old policy $\mathbb{E}_{\begin{subarray}{c}q\sim P(Q)\\ \mathcal{G}_{q^{*}}\sim\pi_{\theta_{\text{old}}^{(s)}}\end{subarray}}$. Consequently, it can be treated as a fixed quantity during differentiation. The gradients $\nabla\rho_{i,t}^{(s)}$ and $\nabla\bar{\rho}_{i,t}^{(s)}$ can be further decomposeds:
@@
Figure 5: Relationship Between the Sign of Advantage and the Direction of Change in Trajectory Output Probabilitys.
@@ VII Conclusion
This paper addresses critical limitations in current reinforcement learning based automated theorem proving (ATP) methods concerning training efficiency and optimization direction by proposing the GDEPO framework. GDEPO systematically enhances the learning efficacy of LLMs in formal reasoning through three key mechanisms. First, it leverages the Lean verifier to precisely identify and discard entirely invalid training samples while actively exploring potentially correct proof paths via a supplementary sampling strategy. Second, it introduces the equal-right advantage mechanism, which preserves the binary correctness signal from the verifier as the primary optimization direction while integrating multidimensional auxiliary rewards to enrich the policy gradient representation. Third, it applies dynamic-iteration backpropagation on challenging samples that initially fail but eventually succeed, accelerating the model’s mastery of difficult yet solvable problems.
@@
This work not only delivers a more efficient and robust training framework for LLM-driven ATP systems but also offers a novel perspective on designing reinforcement learning algorithms for tasks equipped with precise verification mechanisms. With the advancement of AI, nearly all collectible training data have already been exhausted, and the difficulty of acquiring or generating high-value training data continues to rise. While prior research has largely focused on increasing data quantity and quality and on crafting sophisticated reward functions, we argue that future efforts should place equal emphasis on maximizing the utility of existing data and on developing domain-informed strategies for computing trajectory advantages in reinforcement learning applications.

### v2

# GDEPO: Group Dual-dynamic and Equal-right Advantage Policy Optimization with Enhanced Training Data Utilization for Sample-Constrained Reinforcement Learning

@@ Abstract
Automated Theorem Proving (ATP) represents a fundamental challenge in Artificial Intelligence (AI), requiring the construction of machine-verifiable proofs in formal languages such as Lean to evaluate AI reasoning capabilities. Reinforcement Learning (RL), particularly the high-performance Group Relative Policy Optimization (GRPO) algorithm, has emerged as a mainstream approach for this task. However, in ATP scenarios, GRPO faces two critical issues: when composite rewards are used, its relative advantage estimation may conflict with the binary feedback from the formal verifier; meanwhile, its static sampling strategy may discard entire batches of data if no valid proof is found, resulting in zero contribution to model updates and significant data waste. To address these limitations, we propose Group Dual-dynamic and Equal-right Advantage Policy Optimization (GDEPO), a method incorporating three core mechanisms: 1) dynamic additional sampling, which resamples invalid batches until a valid proof is discovered; 2) equal-right advantage, decoupling the sign of the advantage function (based on correctness) from its magnitude (modulated by auxiliary rewards) to ensure stable and correct policy updates; and 3) dynamic additional iterations, applying extra gradient steps to initially failed but eventually successful samples to accelerate learning on challenging cases. Experiments conducted on three datasets of varying difficulty (MinF2F-test, MathOlympiadBench, PutnamBench) confirm the effectiveness of GDEPO, while ablation studies validate the necessity of its synergistic components. The proposed method enhances data utilization and optimization efficiency, offering a novel training paradigm for ATP.
@@ I Introduction
The rapid advancement of Large Language Models (LLMs) has opened new avenues for applying Artificial Intelligence (AI) to formal reasoning. Automated Theorem Proving (ATP), a core challenge in AI, requires systems to construct machine-verifiable mathematical proofs in formal languages such as Lean [3, 4, 31]. Recent progress in ATP on complex theorem-proving tasks stems from the strong prior knowledge and generative capabilities of LLMs, coupled with the availability of extensive training and validation datasets [34, 14, 24, 5, 9, 26]. Notably, AlphaGeometry [23] has demonstrated problem-solving performance approaching International Mathematical Olympiad (IMO) medalist levels, signaling a transition of LLM-driven ATP from theoretical exploration toward practical deployment.
@@
To address the aforementioned issues, we propose the Group Dual-dynamic and Equal-right Advantage Policy Optimization (GDEPO) method, whose workflow is illustrated in Fig. 1. This approach identifies and discards ineffective training data, then re-explores valid solutions. It ensures correct optimization direction under arbitrary reward weight combinations through the equal-right advantage mechanism. Furthermore, for initially invalid training instances that yield correct proofs after re-exploration, the method applies a dynamic number of additional backward passes to rapidly assimilate knowledge that is challenging yet solvable for the model.
@@ II-B Inadequate Adaptability of Policy Optimization Algorithms in Formal Theorem Proving
Inspired by informal LLMs in natural language processing, recent foundational work in formal theorem proving has adopted similar RL-based policy optimization strategies. Current formal theorem provers primarily rely on policy optimization algorithms originally developed for informal LLM training, such as Proximal Policy Optimization (PPO) [21], Direct Preference Optimization (DPO) [17], and Group Relative Policy Optimization (GRPO) [22], which is widely employed by state-of-the-art (SOTA) systems [19, 25, 14]. While these algorithms have demonstrated effectiveness in their native domains—such as language alignment and complex reasoning—they exhibit significant adaptation limitations when transferred to the formal theorem proving setting, where feedback is precise and binary.
@@ II-B3 Limitations of the Group Relative Policy Optimization Paradigm
GRPO has become the core optimization algorithm in current SOTA provers—such as DeepSeek-Prover-V2 [19], Kimina-Prover [25], and Goedel-Prover-V2 [14]—due to its elimination of a separate reward model and computational efficiency. However, its fundamental mechanism, which relies on relative comparisons among samples within a group, inherently conflicts with the absolute correctness criterion demanded by formal verification. This tension manifests in two critical issues. First, when composite rewards are employed, GRPO’s advantage computation may assign negative advantages to otherwise correct proof trajectories due to suboptimal scores in other reward components—a mathematical possibility rigorously established in Section VI. To avoid this, practitioners often resort to binary rewards alone, thereby forfeiting the core capability of RL to align large language model outputs with nuanced human preferences, such as proof conciseness or elegance. Second, GRPO exhibits low data utilization efficiency: in batches containing a high proportion of challenging samples, using a single binary reward frequently results in zero advantages for all trajectories within a group. Consequently, all sampled attempts that fail to produce a valid proof are collectively discarded, wasting valuable training data. To mitigate these risks, existing SOTA systems [19, 25, 14] adopt a conservative strategy—using only binary rewards and deliberately forgoing the potential benefits of composite rewards. Moreover, data batches comprising exclusively correct or incorrect samples are often discarded outright [32]. These compromises yield dual adverse consequences: (i) the precise and rich feedback from formal verifiers remains underutilized, and (ii) the already scarce high-quality training data suffers from further reduced effective usage, impeding the model’s ability to learn from challenging examples. Additionally, the use of a fixed number of optimization iterations treats all samples uniformly, ignoring inherent differences in their learning value. As a result, the current training framework leaves considerable room for more efficient exploitation of formal proof data.
@@
The overall architecture of the GDEPO framework and its comparison with conventional paradigms are illustrated in Fig. 2. The informal large language model RL paradigm, depicted in Fig. 2a, relies on a trained reward model to generate scalar rewards. However, such rewards lack the precise binary signal required to guarantee formal correctness. In contrast, the standard formal ATP paradigm, shown in Fig. 2b, employs a Lean verifier as the ground-truth judge but inherits two rigid design choices from generic RL: (i) a single-round sampling mechanism that struggles to explore low-probability correct solutions for complex problems, and (ii) a fixed number of iterations applied uniformly across all training samples, disregarding their varying learning utility. Moreover, the advantage estimation in this paradigm faces a dilemma: using a composite reward function often leads to misaligned gradient directions, whereas relying solely on the binary correctness reward results in poor data efficiency and weakens the ability of RL optimization to reflect desired output preferences.
@@ III Group Dual-dynamic and Equal-Right Advantage Policy Optimization
GDEPO was proposed to address the aforementioned limitations of standard RL in formal theorem proving. The framework simultaneously improves data efficiency and optimization rationality.
@@ III-B Equal-Right Advantage
The detailed code pipeline is summarized in Algorithm 2. Equal-right advantage decouples the supervision signal: the sign of the advantage—i.e., the direction of policy update—is determined solely by the binary label from the verifier, while auxiliary rewards modulate only the magnitude of the advantage—i.e., the update strength. This mechanism guarantees that all correct proofs receive positive signals and all incorrect ones receive negative signals, preserving the authority of verification outcomes while enabling multi-objective reward shaping. Importantly, it yields correct policy update directions even when all $G$ trajectories for a query are either entirely correct or entirely incorrect.
@@ III-C Dynamic Additional Iterations
Although the dynamic additional sampling module enhances the exploration of low-probability correct solutions, it introduces a new issue of sample heterogeneity: proofs obtained directly in the initial sampling and those discovered only after repeated exploration possess markedly different instructional values for model optimization. The latter—requiring sustained search to succeed—deserve deeper RL attention. Applying identical single-step gradient updates to both cases fails to exploit the richer learning potential of the latter.
To address this, dynamic additional iterations are proposed to allocate computational resources proportionally to each sample’s learning value. The detailed code pipeline is summarized in Algorithm 3. The principle is straightforward: easy queries (correct on the first attempt) and persistently unsolvable hard queries undergo only one update iteration, whereas medium-difficulty queries—those that yield a correct proof only after multiple sampling rounds—are subjected to $m$ iterations to accelerate optimization. Formally, for a query $q$, if its valid output set $\mathcal{G}_{q}^{*}$ contains at least one correct trajectory and was not produced in the first sampling round, then all $G$ trajectories for $q$ are used in $m$ consecutive policy updates. The overall objective function is expressed as:
@@ IV-A Dataset for Training
The experiments are conducted on the publicly available Lean 4 [4] formalized mathematical theorem dataset. To rigorously evaluate GDEPO’s core capability—learning from challenging samples—we construct a targeted high-difficulty training set. The data are sourced from FineLeanCorpus [16], a high-quality dataset comprising formally verified propositions with broad domain coverage and inherently high difficulty. To further isolate highly challenging instances, we use the pass@32 performance of the DeepSeek-Prover-V2-671B model as a filtering criterion, retaining only those samples for which no correct proof is found across 32 independent sampling attempts.
@@ IV-C Baselines
We compare GDEPO-Prover with three SOTA models representing leading paradigms in the field of ATP, ensuring a diverse evaluation across model architectures and optimization strategies.
@@ IV-C1 DeepSeek-Prover-V2
DeepSeek-Prover-V2 [19] adopts a recursive subgoal decomposition paradigm. Its training process integrates multiple stages: problem decomposition and formalization using DeepSeek-V3; recursive subgoal solving with a 7 billion parameter model; curriculum learning for task generation; RL via GRPO algorithm, complemented by supervised fine-tuning and knowledge distillation.
@@ IV-C2 Kimina-Prover
Kimina-Prover [25] emphasizes training stability through KL-constrained RL. It initializes via few-shot fine-tuning, featuring innovations such as automated formalization methods for dataset construction, specialized tokenized reasoning formats for aligning informal and formal reasoning, and techniques like format filtering and negative gradient dropping to enhance stability.
@@ IV-C3 Goedel-Prover-V2
Goedel-Prover-V2 [14] is centered around scaffolded data synthesis and iterative self-correction. Key components include scaffolded data synthesis generating progressively challenging tasks, verifier-guided self-correction leveraging Lean compiler feedback for iterative proof correction, model averaging to mitigate output diversity decline in later training stages by blending model checkpoints, and hybrid GRPO RL incorporating dynamic sampling [32], combined with supervised fine-tuning for efficient theorem proving with smaller parameter sizes.
These benchmark models cover three core technical directions in ATP—modular decomposition, constraint optimization, and feedback-loop enabled synthetic data. By comparing GDEPO, which focuses on innovative training algorithms, against these models, we can precisely isolate the performance gains attributed to our proposed dynamic sampling and advantage estimation mechanisms.
@@ IV-D Training Details
To ensure a fair comparison, all shared hyperparameters of the trainable models are standardized and uniformly configured. The GDEPO algorithm adopts Godel-Prover-V2-8B [14] as the base model and is enhanced within the Open-R1 algorithmic framework [10].
@@ V-A Overall Performance Evaluation
To evaluate the overall performance of GDEPO, we trained GOEDEL-Prover-V2 on a dataset of 30,000 samples containing 30% challenging instances and selected the best-performing model as the representative of our method. Table I reports the performance of this model alongside SOTA baselines, including DeepSeek-Prover-V2 [19], Kimina-Prover [25], and Goedel-Prover-V2 [14], along with their variants of different parameter scales. To ensure that the performance improvement of GDEPO is not affected by uncertainties, the performance data of untrained SOTA models are obtained from their respective official teams. The results show that GDEPO consistently improves performance across all three benchmarks, with a notable 84% gain on the most challenging PutnamBench. This demonstrates that GDEPO can effectively unlock additional performance potential even in models that have approached their performance ceilings, highlighting its strong capability in learning from challenging samples.
TABLE I: Performance Comparison of GDEPO-Prover with Existing SOTA Models and Their Variants Across Model Sizes.
@@
V-B Utilization Effect of Challenging Training Data Evaluation
@@ V-D Ablation Evaluation
TABLE III: Ablation study results, showing the impact of each component on overall performance.
@@ VI-A Advantage Sign as Determinant of Trajectory Probability Direction
Here, $\pi_{\theta_{\text{old}}^{(s)}}(o_{i,t}^{*}\mid q,o_{i,<t}^{*})$ denotes the probability under the old policy and remains constant with respect to $\theta$. The expectation operator $\mathbb{E}$ is representing expectation over the question distribution and the effective response groups sampled from the old policy $\mathbb{E}_{\begin{subarray}{c}q\sim P(Q)\\ \mathcal{G}_{q^{*}}\sim\pi_{\theta_{\text{old}}^{(s)}}\end{subarray}}$. Consequently, it can be treated as a fixed quantity during differentiation. The gradients $\nabla\rho_{i,t}^{(s)}$ and $\nabla\bar{\rho}_{i,t}^{(s)}$ can be further decomposed:
@@
Figure 5: Relationship Between the Sign of Advantage and the Direction of Change in Trajectory Output Probabilities.
@@ VII Conclusion
This paper addresses critical limitations in current RL based ATP methods concerning training efficiency and optimization direction by proposing the GDEPO framework. GDEPO systematically enhances the learning efficacy of LLMs in formal reasoning through three key mechanisms. First, it leverages the Lean verifier to precisely identify and discard entirely invalid training samples while actively exploring potentially correct proof paths via a supplementary sampling strategy. Second, it introduces the equal-right advantage mechanism, which preserves the binary correctness signal from the verifier as the primary optimization direction while integrating multidimensional auxiliary rewards to enrich the policy gradient representation. Third, it applies dynamic-iteration backpropagation on challenging samples that initially fail but eventually succeed, accelerating the model’s mastery of difficult yet solvable problems.
@@
This work not only delivers a more efficient and robust training framework for LLM-driven ATP systems but also offers a novel perspective on designing RL algorithms for tasks equipped with precise verification mechanisms. With the advancement of AI, nearly all collectible training data have already been exhausted, and the difficulty of acquiring or generating high-value training data continues to rise. While prior research has largely focused on increasing data quantity and quality and on crafting sophisticated reward functions, we argue that future efforts should place equal emphasis on maximizing the utility of existing data and on developing domain-informed strategies for computing trajectory advantages in RL applications.